# MiLQ: Benchmarking IR Models for Bilingual Web Search with Mixed Language Queries

### v1

@@ Abstract
Despite bilingual speakers frequently using mixed-language queries in web searches, Information Retrieval (IR) research on them remains scarce. To address this, we introduce MiLQ, Mixed-Language Query test set, the first public benchmark of mixed-language queries, confirmed as realistic and highly preferred. Experiments show that multilingual IR models perform moderately on MiLQ and inconsistently across native, English, and mixed-language queries, also suggesting code-switched training data’s potential for robust IR models handling such queries. Meanwhile, intentional English mixing in queries proves an effective strategy for bilinguals searching English documents, which our analysis attributes to enhanced token matching compared to native queries.
@@
Jonghwi Kim1, Deokhyung Kang1,, Seonjeong Hwang1, Yunsu Kim3, Jungseul Ok1,2, Gary Geunbae Lee1,2, 1 Graduate School of Artificial Intelligence, POSTECH, Republic of Korea , 2 Department of Computer Science and Engineering, POSTECH, Republic of Korea, 3 aiXplain Inc., Los Gatos, CA, USA, {jonghwi.kim, deokhk, seonjeongh, jungseul.ok, gblee}@postech.ac.kr, yunsu.kim@aixplain.com
@@ 1 Introduction
  • We introduce MiLQ, the first public benchmark of mixed-language queries, qualified as realistic and highly preferred by bilinguals.
@@
SO 151 16.2$\rightarrow$59.6 2.38 2.34 2.73 2.58 2.95 5.4$\rightarrow$36.1 2.76 2.34 2.63 2.51 2.77
FI 151 7.3$\rightarrow$40.2 2.48 2.52 2.79 2.70 2.59 2.2$\rightarrow$45.3 2.63 2.15 2.63 2.44 2.28
DE 151 9.1$\rightarrow$61.8 2.67 2.68 2.61 2.50 2.21 2.1$\rightarrow$41.1 2.55 2.11 2.43 2.15 1.80
FR 151 5.7$\rightarrow$35.0 2.52 2.55 2.84 2.51 2.31 2.3$\rightarrow$32.9 2.80 2.30 2.84 2.51 2.31
@@
FA 45 2.2$\rightarrow$15.0 2.98 2.98 2.87 2.82 2.64 0.1$\rightarrow$5.6 3.00 2.93 2.91 2.81 2.68
RU 44 0.0$\rightarrow$51.7 2.89 2.50 2.72 2.30 2.16 0.6$\rightarrow$51.7 2.93 2.45 2.73 2.16 2.14
Average 111.4 6.2$\rightarrow$39.5 2.64 2.60 2.78 2.59 2.57 5.0$\rightarrow$31.6 2.80 2.45 2.76 2.45 2.46
Table 1: Quality measurements for MiLQ (Title & Description queries). Code-Mixing Index (CMI) is on a 0-100 scale (Original Query CMI $\rightarrow$ Mixed-language Query (MiLQ) CMI). For GPT-Eval (Accuracy [Acc.] & Fluency [Flu.]) and Human-Evaluation (Acc. & Flu. & Realism [Real.]), both on a 1-3 scale, cell backgrounds are colored in a fine-grained red gradient from lightest red (scores $\approx$1.0) to darkest red (scores $\approx$3.0). The ’Average’ row is bolded. "XX" denotes the native language.
@@ 2 MiLQ: Mixed-Language Query test set
We measured MiLQ’s quality considering its language mixing, meaning preservation, naturalness, and realism (Table 1). First, for language mixing, we used the Code-Mixing Index (CMI) Das and Gambäck (2014) (0-100 scale, higher=more mixing; Appendix A.3). Average CMI increased from 6.2 to 39.5 (Title) and 5.0 to 31.6 (Description), showing substantially more mixing than originals. Next, GPT-Eval (GPT-4o) using Kuwanto et al.’s framework (high human alignment, Kendall’s Tau > 0.5) assessed MiLQ (1-3 scale; rubrics in Appendix A.4) for Accuracy (Acc.) (meaning preservation, correct term use) and Fluency (Flu.) (naturalness, readability, seamlessness). MiLQ achieved strong average GPT-Eval scores: Acc. 2.64 / Flu. 2.60 (Title) and Acc. 2.80 / Flu. 2.45 (Description). Lastly, for Human-Eval, three bilingual annotators per query assessed MiLQ on a 1-3 scale (detailed guidelines in Appendix A.5). This evaluation covered Accuracy (Acc.) and Fluency (Flu.), using criteria consistent with GPT-Eval, and an additional Realism (Real.). Realism specifically assessed how naturally bilingual speakers might use the given mixed-language query in real search scenarios. Human evaluators rated MiLQ highly, with average scores: Acc. 2.78 / Flu. 2.59 / Real. 2.57 (Title) and Acc. 2.76 / Flu. 2.45 / Real. 2.46 (Description). These consistently high scores in all metrics affirm the quality and reliability of MiLQ.
@@ 3 Experimental Setup
Test Scenarios & Collections
@@ 4 Results and Analysis
Figure 3: Token-level similarity matrices from Cross-R for German and mixed-language queries on ground truth passage. The y-axis shows tokenized queries (mixed-language left, native right), and the x-axis represents the tokenized English passage. MaxSim tokens are marked by $\times$, and the code-switched parts are highlighted.
@@
Our analysis compared MaxSim token pair similarity (a query token and its maximal similarity document token) in mixed-language versus native queries. Figure 3 (left) shows mixed-language queries, by including English terms (e.g., "Intellectual Property Rights" from German "Intellektuelle Eigentumsrechte"), allow these English tokens to form MaxSim pairings ($\times$) with accurate, higher similarity scores. Conversely, native queries (right) rely on cross-lingual interpretation of native tokens (e.g., German "Intellekt," "Eigen") to map English concepts. While MaxSim pairings are also identified ($\times$), this mapping yields weaker similarity for such crucial English concepts. Thus, intentionally mixing English terms improves MaxSim matching through higher similarity scores for English terms—a key rationale (RQ3) for MQIR’s enhanced English retrieval.
@@ 5 Conclusion
This study addressed prevalent yet understudied mixed-language querying by bilinguals by introducing MiLQ—the first public user-crafted MQIR benchmark, confirmed for realism and high user preference. Experiments on MiLQ revealed inconsistent performance of current IR models on diverse queries, underscoring the need for robust models and the potential of code-switched training data. Finally, we find that intentional English mixing boosts English document retrieval for bilinguals.
@@ 6 Limitations
In our research, we use NeuCLIR22 (Lawrie et al., 2023a), which are licensed under CC BY-NC 4.0. The models GPT-4o (Hurst et al., 2024) are licensed under OpenAI, Apache-2.0, All models were used strictly for research purposes, and no artifacts were utilized beyond the scope of the study.
@@ A.6 Insights from Annotator Interviews on Mixed-Language Query Usage
To gain deeper insights into why and when bilingual users employ mixed-language queries (MiLQ) in real-world online searches, we conducted semi-structured interviews with all annotators. A common question posed was: "In what situations are mixed-language queries commonly used in real-world online search contexts, and for what reasons?" Table 4 summarizes the key themes derived from their responses.
@@ B.1 Benchmark Statistics
ZH FA RU SW SO FI DE FR
# Queries 47 45 44 151 151 151 151 151
# Documents 3.2M 2.2M 4.6M 113K 113K 113K 113K 113K
# Passages 19.8M 14.0M 25.1M 1.02M 1.02M 1.02M 1.02M 1.02M
@@ B.3 Implementation Details
Our retrieval experiments are based on the ColBERT architecture Khattab and Zaharia (2020), a multi-vector approach for dense retrieval. We utilized the publicly available PLAID-X implementation [^6] https://github.com/hltcoe/ColBERT-X for all model training and inference. Consistent with standard ColBERT practices, most training artifacts and hyperparameters were adopted directly. Our primary modification involved setting the maximum document passage length to 180 tokens. Following established methods Bendersky and Kurland (2008); Dai and Callan (2019), documents longer than this threshold were segmented into 180-token passages. During evaluation, the score for each document was determined using the maximum passage score (MaxP) strategy.
@@ B.4 Performance in Individual Languages
Query Lang BM25 Mono-Distil Mixed-Distil Cross-Distil mContriever NMT $\rightarrow$ BM25 NMT $\rightarrow$ Mono-Distil

### v2

@@ Abstract
Despite bilingual speakers frequently using mixed-language queries in web searches, Information Retrieval (IR) research on them remains scarce. To address this, we introduce MiLQ, Mixed-Language Query test set, the first public benchmark of mixed-language queries, qualified as realistic and relatively preferred. Experiments show that multilingual IR models perform moderately on MiLQ and inconsistently across native, English, and mixed-language queries, also suggesting code-switched training data’s potential for robust IR models handling such queries. Meanwhile, intentional English mixing in queries proves an effective strategy for bilinguals searching English documents, which our analysis attributes to enhanced token matching compared to native queries. [^1] The code for this work are available at : https://github.com/jonghwi-kim/milq
@@
Jonghwi Kim1, Deokhyung Kang1, Seonjeong Hwang1, Yunsu Kim3thanks: This work was done when the author was at aiXplain, Jungseul Ok1,2, Gary Geunbae Lee1,2, 1 Graduate School of Artificial Intelligence, POSTECH, Republic of Korea , 2 Department of Computer Science and Engineering, POSTECH, Republic of Korea, 3 Lilt, Inc. {jonghwi.kim, deokhk, seonjeongh, jungseul.ok, gblee}@postech.ac.kr, yunsu.kim@lilt.com
@@ 1 Introduction
  • We introduce MiLQ [^3] https://catalogue.elra.info/en-us/repository/browse/ELRA-E0047/ , the first public benchmark of mixed-language queries, qualified as realistic and relatively preferred by bilinguals.
@@
[…]
Table 1: Quality measurements for MiLQ (Title & Description queries). Code-Mixing Index (CMI) is on a 0-100 scale (Original Query CMI $\rightarrow$ Mixed-language Query (MiLQ) CMI). For GPT-Eval (Accuracy [Acc.] & Fluency [Flu.]) and Human-Evaluation (Acc. & Flu. & Realism [Real.]), both on a 1-3 scale, cell backgrounds are colored in a fine-grained red gradient from lightest red (scores $\approx$1.0) to darkest red (scores $\approx$3.0). The ’Average’ row is bolded. ”XX” denotes the native language.
@@ 2 MiLQ: Mixed-Language Query test set
We measured MiLQ’s quality considering its language mixing, meaning preservation, naturalness, and realism (Table 1). First, for language mixing, we used the Code-Mixing Index (CMI) Das and Gambäck (2014) (0-100 scale, higher=more mixing; Appendix A.3). Average CMI increased from 6.2 to 39.5 (Title) and 5.0 to 31.6 (Description), showing substantially more mixing than originals. Next, GPT-Eval (GPT-4o) using Kuwanto et al.’s framework (high human alignment, Kendall’s Tau ¿ 0.5) assessed MiLQ (1-3 scale; rubrics in Appendix A.4) for Accuracy (Acc.) (meaning preservation, correct term use) and Fluency (Flu.) (naturalness, readability, seamlessness). MiLQ achieved strong average GPT-Eval scores: Acc. 2.64 / Flu. 2.60 (Title) and Acc. 2.80 / Flu. 2.45 (Description). Lastly, for Human-Eval, three bilingual annotators per query assessed MiLQ on a 1-3 scale (detailed guidelines in Appendix A.5). This evaluation covered Accuracy (Acc.) and Fluency (Flu.), using criteria consistent with GPT-Eval, and an additional Realism (Real.). Realism specifically assessed how naturally bilingual speakers might use the given mixed-language query in real search scenarios. Human evaluators rated MiLQ highly, with average scores: Acc. 2.78 / Flu. 2.59 / Real. 2.57 (Title) and Acc. 2.76 / Flu. 2.45 / Real. 2.46 (Description). These consistently high scores in all metrics affirm the quality and reliability of MiLQ.
@@ 3 Experimental Setup
Test Scenarios & Dataset
@@ 4 Results and Analysis
Figure 3: Token-level similarity matrices from Cross-Distill for German and mixed-language queries on ground truth passage. The y-axis shows tokenized queries (mixed-language left, native right), and the x-axis represents the tokenized English passage. MaxSim tokens are marked by $\times$, and the code-switched parts are highlighted.
@@
Our analysis compared MaxSim token pair similarity (a query token and its maximal similarity document token) in mixed-language versus native queries. Figure 3 (left) shows mixed-language queries, by including English terms (e.g., ”Intellectual Property Rights” from German ”Intellektuelle Eigentumsrechte”), allow these English tokens to form MaxSim pairings ($\times$) with accurate, higher similarity scores. Conversely, native queries (right) rely on cross-lingual interpretation of native tokens (e.g., German ”Intellekt,” ”Eigen”) to map English concepts. While MaxSim pairings are also identified ($\times$), this mapping yields weaker similarity for such crucial English concepts. Thus, intentionally mixing English terms improves MaxSim matching through higher similarity scores for English terms—a key rationale (RQ3) for MQIR’s enhanced English retrieval.
@@ 5 Conclusion
This study addressed the prevalent yet understudied phenomenon of mixed-language querying among bilingual speakers by introducing MiLQ—the first public user-crafted MQIR benchmark, validated for both realism and high user preference. Our comprehensive experiments on MiLQ revealed that current IR models exhibit inconsistent performance across diverse query types, highlighting the need for more robust retrieval systems and demonstrating the promising potential of code-switched training data. Finally, we discovered that intentionally mixing English terms into queries serves as an effective strategy for enhancing English document retrieval among bilingual users.
@@ 6 Limitations
Dataset Licensing and Usage
Our work uses three primary datasets: NeuCLIR22 Lawrie et al. (2023a), CLEF00-03 Braschler (2003), and our newly introduced MiLQ dataset. We have verified the licensing terms for all existing datasets and ensured our usage is consistent with their intended research purposes. The NeuCLIR22 dataset is published by NIST with public access level and is subject to the NIST Open License. The CLEF00-03 data is distributed by ELDA under an End-User Agreement for Evaluation Packages for Research Use, which permits evaluation purposes. Our MiLQ dataset will be distributed by ELDA under a free evaluation license for academic organizations. The dataset is accessible through our code repository. The MiLQ dataset and our findings are intended solely for academic research purposes in multilingual information retrieval. We discourage commercial deployment without further evaluation of potential societal impacts and biases.
Human Annotation
For our MiLQ dataset, queries were created by bilingual speakers and subsequently validated by a different group of bilingual speakers to ensure quality and reduce bias. All participants in both stages were compensated fairly for their work based on regional wage standards and time estimates for each task. The detailed annotation guidelines and compensation structure for each stage are described in Appendix A.1 and A.5. We obtained informed consent from all participants, clearly explaining how their contributions would be used in research and dataset creation.
Potential Risks
The scope of our study is limited to the news domain and nine languages: English, Swahili, Somali, Finnish, German, French, Chinese, Persian, and Russian. Therefore, our findings may not generalize to other domains, genres, or languages not represented in our evaluation. Our findings may inadvertently favor certain language pairs or retrieval approaches that work better for high-resource languages, potentially contributing to digital language divides. Regarding personal information, we followed the existing privacy protection measures established by NIST and ELDA for the original datasets.
Acknowledgments
This research was supported by the MSIT (Ministry of Science and ICT), Korea, under the ITRC(Information Technology Research Center) support program (IITP-2025-RS-2020-II201789, Contribution Rate: 47.5%) supervised by the IITP (Institute for Information & Communications Technology Planning & Evaluation); by the Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2025 (Project Name: Development of an AI-Based Korean Diagnostic System for Efficient Korean Speaking Learning by Foreigners, Project Number: RS-2025-02413038, Contribution Rate: 47.5%); and by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH), Contribution Rate: 5%).
The MiLQ dataset is released through the ELRA catalogue under the Evaluation Use - ELRA EVALUATION license, providing free access to researchers for evaluation purposes. If you use this dataset, please include the following citation in your acknowledgements: MiLQ: Mixed-Language Query Test Set for Bilingual Web Search - Evaluation Package, ELRA catalogue (http://catalog.elra.info), ISLRN: 317-005-302-361-6, ELRA ID: ELRA-E0047
The following dataset was used for evaluation in this study: The CLEF Test Suite for the CLEF 2000-2003 Campaigns – Evaluation Package, ELRA catalogue (http://catalog.elra.info), ISLRN: 200-586-423-805-2, ELRA ID: ELRA-E0008
@@ A.6 Insights from Annotator Interviews on Mixed-Language Query Usage
To gain deeper insights into why and when bilingual users employ mixed-language queries (MiLQ) in real-world online searches, we conducted semi-structured interviews with all annotators. A common question posed was: ”In what situations are mixed-language queries commonly used in real-world online search contexts, and for what reasons?” Table 4 summarizes the key themes derived from their responses.
@@ B.1 Benchmark Statistics
ZH FA RU SW SO FI DE FR EN
# Queries 47 45 44 151 151 151 151 151 151
# Documents 3.2M 2.2M 4.6M 113k
# Passages 19.8M 14.0M 25.1M 1.01M
@@ B.3 Implementation Details
For BM25, we utilized the Pyserini toolkit Lin et al. (2021), which provides reproducible sparse retrieval through Lucene  [^10] https://github.com/castorini/pyserini. Dense retrieval experiments with mContriever Izacard et al. employed facebook/mcontriever-msmarco model [^11] https://huggingface.co/facebook/mcontriever-msmarco.
Our primary retrieval experiments use the ColBERT architecture Khattab and Zaharia (2020), a multi-vector approach for dense retrieval. We utilized the publicly available PLAID-X implementation [^12] https://github.com/hltcoe/ColBERT-X for all model training and inference. Consistent with standard ColBERT practices, most training artifacts and hyperparameters were adopted directly. Our primary modification involved setting the maximum document passage length to 180 tokens. Following established methods Bendersky and Kurland (2008); Dai and Callan (2019), documents longer than this threshold were segmented into 180-token passages. During evaluation, the score for each document was determined using the maximum passage score (MaxP) strategy.
All experiments were conducted with a single run. Due to the approximate nearest neighbor (ANN) search employed in ColBERT, experimental results may exhibit minor variations depending on the indexing process. However, we observed that such variations do not lead to substantial differences.
@@ B.4 Performance in Individual Languages
[…]